araba and john are partners in a business they zgree to share their profit in the ratio 2:3.if the total annual profit is 6500.how much will each get
Answer by DENNIS DON(8)

Let A represent araba and J represent john.
A's ratio =2
J's ratio =3
Total ratio =2+3=5
A's share =(2/5)*6500 = 2600
J's share =(3/5)*6500 = 3900